# Predictive Reliability Assessment of Distribution Grids with Residential Distributed Energy Resources

### v1

@@ V-C Joint adoption scenarios
V-D Application to RBTS Bus 4
The RBTS test system is made up of 7 feeders, 38 load points, and 4779 customers, of which 22 load points serve a total of 4700 residential customers [22]. The test system is modified as follows: Feeders 2,5, and 6, which serve nine industrial customers, are removed. Since the granularity of the simulation is at the first abstraction layer, the necessary condition is to have uniform reliability characteristics for all the load points across the distribution system. To achieve this, the system indices for the RBTS system, excluding the industrial feeders, are calculated. We use these indices as the basic load point indices for all the load points in the modified system. In this setup, we consider that commercial customers do not have any onsite DER. Therefore, the commercial customers’ AIF and AID will equal the load point failure and repair rate.
We chose the location of Los Angeles County for the simulation and calculated the representative residential load profile for the location from the ResStock dataset [23]. We used the solar irradiance data from [24], at the chosen location to calculate the rooftop PV system output in the simulation. We performed the MCS for the 16 joint adoption scenarios in Table II and tabulated the SAIFI and SAIDI values from the simulation results in Tables III and IV.
@@ V-D Application to RBTS Bus 4
Limited 0.2978 0.2982 0.29645 0.2962
Varied 0.2836 0.1894 0.2007 0.1317
Median 0.2850 0.1853 0.1759 0.0868
Highly Conc. 0.2738 0.1535 0.1463 0.0183
@@
Limited 2.965 2.942 2.933 2.932
Varied 1.981 1.351 1.305 0.967
Median 1.884 1.172 0.954 0.511
Highly Conc. 1.729 0.923 0.703 0.095
Figure 5: Density plots on the left represent the Joint adoption density of PV and ES for the corresponding scenarios on the right. Density plots on the right represent the variation in AIF for the annotated scenarios.
The simulation results show the expected outcomes with SAIFI and SAIDI. As the adoption of PV and ES grows the system reliability metrics improve. Fig. 5, visualizes two selected adoption scenarios and the respective AIF index plots. The first set of subplots on the top correspond to the {Limited, Median-Focused} adoption scenario. The joint density contour on the left shows the PV and ES adoption concentration. The plot shows that the PV adoption is shallow between the range of having no PV to a PV size of one times the peak load. Meanwhile, the concentration of ES adoption is around the median value of 3. The corresponding density plot on the right shows the density of the AIF index constructed from the AIF values of the samples drawn from joint density. When we look at the estimated SAIFI value, we see a slight improvement from the baseline, but this does not tell the entire story. The variation in AIF, even though small, shows us that for some configurations of PV and ES sizes, the AIF deteriorates compared to the baseline. The following conditions explain the deterioration in the AIF values: if there is a long enough grid interruption, the end user with small PV and ES sizes will experience multiple interruptions due to insufficient power from the DER system and the cyclic nature of the PV generation. This phenomenon is reported in the literature from studies performed on a single residential unit. Our method quantifies and captures these intricate details at the residential level for an entire distribution system, which would otherwise be lost when a single metric is used to quantify system reliability.
The second set of subplots shows the {Varied, varied} adoption scenario. The joint density contours on the left show that the PV and ES adoption spans the entire range, and there is a high positive correlation. The corresponding density plot shows a wide variation in the AIF index. As with the previous scenario, the estimated AIF shows huge improvement, but the single numerical metric does not capture the entire variation in the AIF index. There are two peaks in the distribution, and the estimated SAIFI falls in between the peaks. The peak on the right of the estimated SAIFI has a higher density than the peak on the left, indicating that customers with higher PV and ES sizes contribute the most to the improvement in the estimated SAIFI. A single numerical metric quantifying the estimated SAIFI will not capture these variation and cannot provide detailed residential-level insights.While we simulated sixteen PV and ES joint scenarios, we explained only two scenarios due to space limitations. The rest of the scenarios are visualized in Figs. 5(a) and 5(b).
(a) SAIFI distributions for various adoption scenarios.
(b) SAIDI distributions for various adoption scenarios.
Figure 6: Each subplot is for a specific PV adoption and each distribution is for a specific ES adoption as indicated in the legend.
@@ VI Conclusion
This work proposes a bottom-up probabilistic approach to integrate residential DER into reliability evaluation. The variability of the residential DERs is captured and integrated into the assessment using micro-level penetration modeling. Applying this approach to the RBTS test system shows that a single metric, such as SAIFI or SAIDI, will not capture the variation in reliability experienced by the end users in the presence of BTM DERs. The proposed approach provides a powerful way to quantify the variability in the reliability indices. The findings of this research are particularly relevant for utilities and policymakers, providing them with a robust framework to anticipate and manage the challenges posed by the rapid integration of residential DERs. Future work will further evaluate the impact of adopting this probabilistic approach in speeding up the transition towards more renewable energy resources in large-scale testing.

### v2

@@ IV-D Joint adoption scenarios
IV-E Modeling Assumptions and Applicability
To clarify the scope of the proposed framework, we summarize below the key assumptions adopted in this work and discuss their applicability across different distribution system contexts.
Uniform Load Point Reliability: In the first abstraction layer, it is assumed that the reliability indices are uniform at all load points. This assumption is appropriate for urban and suburban systems, where infrastructure quality and operational standards are relatively homogeneous. However, in rural networks, the length of the feeder and the customer density often result in a heterogeneous reliability performance. The proposed framework is designed to accommodate such nonuniform load point indices through a second abstraction layer. In this paper, we focus on the uniform case to illustrate the methodology; extending the analysis to heterogeneous load point reliability will be pursued in future work.
Homogeneity of Normalized Load Profiles: Residential load shapes are assumed to be homogeneous once normalized by peak demand. Empirical datasets such as NREL’s ResStock [21] confirm that load profiles within a climatic and socio-economic region exhibit consistent temporal patterns, with variation largely explained by magnitude rather than shape.
PV and ES Component Reliability: The framework also accounts for the reliability of behind-the-meter DER components. Rooftop PV is modeled through AC modules (PV panel plus micro-inverter), and storage through inverter-integrated battery modules, each subject to independent failure and repair processes. Including these subsystem dynamics ensures that customer-level reliability indices reflect not only grid-side interruptions but also the availability of the DER hardware itself. This treatment extends the framework beyond adoption patterns alone and aligns with prior component-level reliability models established in our earlier work [19].
Adoption Distributions: Marginal adoption distributions for PV and ES are modeled using theoretical forms (Beta, Uniform, Truncated Normal) to represent different adoption archetypes. This scenario-based approach is consistent with prior distributed energy studies [20, 22, 23]. The objective is not to forecast specific adoption trajectories, but to illustrate how the framework evaluates system reliability under diverse but plausible penetration patterns.
PV–ES Correlation: Correlation coefficients in the joint adoption scenarios (Table II) are selected to span a plausible range of interactions. While detailed empirical correlation data remain limited, emerging evidence suggests strong complementarity. For example, the NREL LA100 study [24] projects 1.1–1.5 GW of distributed storage co-adopted with 3–4 GW of rooftop PV in Los Angeles by 2045, corresponding to roughly 30–40% of PV capacity being paired with storage. In the near term, EIA data [25] indicate that more than 50% of new residential PV installations in California in 2024 were paired with storage. Together, these findings motivate the inclusion of moderate-to-high positive correlations (0.3–0.8) in the scenario matrix. At the same time, survey and behavioral studies [26, 27, 28] show that adoption decisions are also shaped by user perception, economic considerations, and policy incentives, meaning that weaker correlations (0.1–0.3) may also arise in practice. Finally, a small negative correlation (-0.2) is included as a stress-test case to explore conditions where storage adoption may be incentivized independently of PV (e.g., as a resilience investment for backup-only applications). Overall, the correlation scenarios are intended to provide sensitivity bounds rather than precise forecasts, ensuring the robustness of the proposed reliability assessment framework.
As empirical adoption datasets become available at higher granularity, the proposed framework is designed to directly integrate such data, ensuring that the current scenario-based analysis can evolve into a data-driven validation
V Application of the probabilistic approach
V-A Monte Carlo Simulation Framework
The reliability indices under different DER adoption scenarios are estimated using an adaptive Monte Carlo (MC) simulation framework. The process is structured as a sequence of independent stratified sampling batches, each covering the input probability space with low discrepancy. Within each batch, the system model is evaluated and per-sample reliability metrics (e.g., SAIFI, SAIDI) are accumulated. The overall workflow, shown in Fig. 4, iteratively expands the sample size until statistical convergence is achieved. This adaptive strategy avoids the inefficiencies of fixed sample sizes and ensures that precision requirements are met without unnecessary computation.
Estimator updates. For each reliability metric $k\in\{\text{SAIFI, SAIDI}\}$, running means $\hat{\mu}_{k}(n)$ and unbiased variances $\hat{\sigma}^{2}_{k}(n)$ are updated in streaming form (Welford’s method). This guarantees numerical stability and requires only $O(1)$ memory.
Stopping rule. Convergence is determined using two criteria:
  1. 1.
    Minimum information floor: At least one full batch ($n_{\min}=10$) must be accumulated before any inference is allowed.
  2. 2.
    Confidence interval precision: For each metric, the half-width
    $h_{k}(n)=z_{\alpha/2}\,\hat{\sigma}_{k}(n)/\sqrt{n}$ (43)
    must not exceed a predefined tolerance. In this study, tolerances are set as
    $\varepsilon_{\text{SAIFI}}=0.005~\text{interruptions/customer-year},$
    $\varepsilon_{\text{SAIDI}}=0.1~\text{hours/customer-year}.$
    These thresholds correspond to reporting precision at the second decimal place for SAIFI and the first decimal place for SAIDI, consistent with IEEE Std. 1366 reporting practices.
This criterion ensures that the reported estimates have a reproducible precision level and prevents premature termination due to insufficient sampling.
Parameterization in this study. We adopt a 95% confidence level ($\alpha=0.05$), batch size of 10 (replicate-stratified mode), a maximum sample cap of 2000, and the absolute CI half-width thresholds listed above. Deterministic seeding (42) and per-batch hashing ensure reproducibility, while the final number of samples and achieved half-widths are reported for transparency.
With these convergence criteria and precision targets established, the following subsection applies the adaptive MC framework to the RBTS Bus 4 system.
Figure 4: MCS flowchart.
V-B Application to RBTS Bus 4
System setup and abstraction: The RBTS distribution test system consists of seven feeders, 38 load points, and 4779 customers [29]. For this study, feeders 2, 5, and 6, which supply nine industrial customers are removed, leaving approximately 4700 residential customers served by 22 load points. Because the analysis in this paper is conducted at the first abstraction layer, all load points are assigned uniform reliability characteristics. To enforce this, the baseline system indices of the modified RBTS (excluding industrial feeders) are computed and applied as the basic load point indices across the network. This homogenization ensures that the computational effort depends primarily on the number of Monte Carlo samples rather than the system size, enabling the framework to scale to larger networks without increasing simulation complexity at this level of abstraction. Commercial customers are assumed not to host any behind-the-meter DER, and their reliability indices therefore equal the underlying load point values.
Inputs and scenarios: The case study uses Los Angeles County as the representative location. Residential load profiles are drawn from the ResStock dataset [21], and rooftop PV generation is computed from the NSRDB irradiance data [30]. Since ResStock internally uses NSRDB irradiance as its solar input, the load and generation data are inherently consistent, eliminating potential mismatches between demand and resource drivers. The adoption distributions for PV and ES, along with their correlation structures, are defined in Section IV-C and summarized in Table II. These distributions are treated as scenario-based abstractions rather than forecasts, enabling the framework to capture a range of plausible adoption patterns.
Simulation results: Tables III and IV summarize the system-wide reliability indices across the sixteen PV–ES adoption scenarios. For reference, the baseline RBTS values without any DER are $\text{SAIFI}=0.30$ f/yr and $\text{SAIDI}=3.47$ h/yr. The simulated results confirm that increased adoption of PV and ES generally improves both indices, but the magnitude of improvement varies substantially across adoption patterns.
In the best-case scenario with both PV and ES highly concentrated ({HC, HC}), the indices reduce to $\text{SAIFI}=0.0211$ f/yr and $\text{SAIDI}=0.107$ h/yr. This corresponds to a reduction of approximately 93% in interruption frequency and 97% in interruption duration relative to the baseline, highlighting the potential value of coordinated adoption. At the other extreme, the limited-adoption case ({L, L}) produces indices nearly identical to the baseline ($\text{SAIFI}=0.3154$, $\text{SAIDI}=3.300$). Interestingly, in some low-adoption cases, the simulated SAIFI is slightly higher than the baseline. This is because the framework accounts not only for grid-side interruptions but also for the stochastic reliability of the PV and ES subsystems themselves. When PV adoption is shallow and ES capacity is small, these resources may deplete quickly or fail to sustain load over the full duration of a utility outage. As a result, a single load point interruption can manifest as multiple end-user interruptions due to PV cycling and insufficient ES support. This phenomenon has been observed in our prior work [19], where small DER sizes led to higher interruption counts even as average interruption duration improved. Intermediate cases, such as the {MF, MF} scenario, yield $\text{SAIFI}=0.1933$ and $\text{SAIDI}=1.070$ h/yr, demonstrating substantial but less dramatic gains.
These results illustrate that the proposed framework can capture both extreme improvements and nuanced degradation mechanisms, and that its estimates are consistent with baseline expectations in low-adoption cases.
To illustrate these findings more concretely, we examine two representative scenarios in detail. The first highlights the maximum potential benefit under coordinated adoption, while the second reveals the heterogeneity of outcomes under diverse adoption patterns.
To illustrate these findings more concretely, we examine two representative scenarios in detail. The first highlights the maximum potential benefit under coordinated adoption, while the second reveals the heterogeneity of outcomes under diverse adoption patterns.
Highly concentrated adoption ({HC, HC}): In the highly concentrated scenario, both PV and ES capacities cluster at the upper end of their respective ranges. This leads to a dramatic improvement in reliability: $\text{SAIFI}\approx 0.021$ f/yr and $\text{SAIDI}\approx 0.107$ h/yr, corresponding to reductions of about 93% and 97% relative to the baseline. The AIF distribution is extremely tight, indicating that nearly all customers experience similar reliability gains. This case demonstrates the maximum potential of coordinated PV–ES adoption and shows how the framework captures significant system-wide improvements when adoption is both high and uniform.
Varied adoption ({V, V}): In contrast, the varied adoption scenario exhibits wide diversity in PV and ES capacities, spanning nearly the entire feasible range. The resulting AIF distribution is clearly bimodal: one cluster of customers benefits significantly, achieving reliability comparable to the {HC, HC} case, while another cluster remains near baseline. The estimated mean SAIFI ($\sim$0.21 f/yr) lies between these peaks, but does not represent either group well. This example highlights the central contribution of the framework: single numerical indices obscure the heterogeneity of customer outcomes, whereas the distributional perspective reveals both the opportunities and the risks of uneven DER adoption.
@@ V-B Application to RBTS Bus 4
Limited 0.3154 0.3150 0.3129 0.3127
Varied 0.3047 0.2053 0.2167 0.1540
Median 0.3018 0.2074 0.1933 0.0987
Highly Conc. 0.2867 0.1714 0.1603 0.0211
@@
Limited 3.300 3.298 3.281 3.279
Varied 2.277 1.591 1.532 1.244
Median 1.993 1.327 1.070 0.577
Highly Conc. 1.813 1.049 0.787 0.107
Figure 5: Joint adoption densities (left) and AIF distributions (right) for selected PV–ES adoption scenarios. The top row corresponds to highly concentrated adoption of PV and ES ({HC, HC}). The bottom row corresponds to varied adoption of PV and ES ({V, V}). The red dashed line marks the baseline SAIFI (0.30 f/yr) and the black dashed line marks the estimated mean SAIFI for each scenario.
TABLE V: Representative adaptive Monte Carlo convergence statistics across PV–ES adoption scenarios. Reported cases correspond to the fastest, median, and slowest scenarios.
Scenario Samples Runtime (s)
Fastest (L, L) 30 27
Median (MF, V) 275 106
Slowest (HC, V) 2000 1520
Computational efficiency and scalability: The adaptive Monte Carlo framework converged in a median of approximately 275 samples, corresponding to a runtime of about 2 minutes on a standard workstation. The fastest scenarios required as few as 30 samples (27 s), while the most challenging scenario reached the prescribed maximum of 2000 samples (about 25 minutes). A summary of representative convergence statistics is provided in Table V. This adaptive strategy therefore avoids the inefficiencies of fixed-sample simulations, achieving precision targets with tractable runtimes across all cases. At the first abstraction layer considered in this paper, the computational cost depends primarily on the number of samples and is independent of the network size, since all load points share uniform baseline reliability. For more detailed studies, a distribution system can be partitioned into a small number of reliability regions, each representing a set of load points with homogeneous reliability characteristics. The computational cost then scales with the number of regions rather than the total number of load points, which preserves tractability even for large feeders. Moreover, the batch-based sampling architecture ensures that the framework remains amenable to parallelization, supporting scalability for realistic distribution system applications.
Uncertainty in input distributions: The reliability estimates reported here are subject to both statistical and distributional uncertainty. Statistical uncertainty is controlled through the adaptive Monte Carlo procedure, which enforces confidence interval half-widths of 0.005 for SAIFI and 0.1 h/yr for SAIDI. Distributional uncertainty arises from the use of scenario-based adoption models and correlation coefficients, which are intended as plausible sensitivity cases rather than forecasts. These results should therefore be interpreted as sensitivity analyses around plausible adoption pathways. Explicit quantification of distributional uncertainty remains an important direction for future work.
@@ VI Conclusion
This paper introduced a bottom-up probabilistic framework for predictive reliability assessment of distribution systems with residential DERs. By modeling adoption at the micro level and explicitly incorporating PV and ES component reliability, the framework captures distributional variability in reliability outcomes rather than relying solely on system-average indices. Application to the modified RBTS Bus 4 system demonstrated that conventional indices such as SAIFI and SAIDI mask significant heterogeneity in customer-level experiences when DERs are present.
Numerical results showed that high DER adoption can reduce SAIFI and SAIDI by more than 90% compared to the baseline, while shallow or uneven adoption can in some cases worsen interruption frequency. These insights highlight the importance of assessing not only mean reliability improvements but also the spread of outcomes across different customer configurations. The findings provide utilities and regulators with a robust way to anticipate the reliability impacts of DER adoption pathways and to design policies that capture their full value.
Future work will extend this framework to larger, more detailed distribution systems by incorporating multiple reliability regions, empirical adoption datasets, and resilience-focused scenarios. These enhancements will enable more comprehensive planning studies that align with the rapid growth of residential solar and storage.